Zeronknight (zeronknight):

Solve the following inequality and then choose the correct solution. -\frac{x}{9}\; -\; 2< 3 A. x> -\frac{5}{9} B. x> -45 C. x< -45 D. x< -\frac{5}{9}

Zeronknight (zeronknight):

\[ -\frac{x}{9}\; -\; 2< 3\]

Zeronknight (zeronknight):

\[A. x> -\frac{5}{9}\]

Zeronknight (zeronknight):

@sleepyjess

OpenStudy (anonymous):

Where is your work?

OpenStudy (sleepyjess):

Inequalities! I'm doing this right now in my class! Let's isolate x

Zeronknight (zeronknight):

I don't even know how to even do this, so I can'y show my work :P

OpenStudy (anonymous):

I'll let Jess take over if she wants to :)

Zeronknight (zeronknight):

Ok, we need to isolate X. What do fractions have to do with inequalities though? They are very confusing :\

OpenStudy (sleepyjess):

zeron, do you know how to solve \(-\dfrac x9 -2 = 3\)

Zeronknight (zeronknight):

Nope

Zeronknight (zeronknight):

I know how to it with no fractions, or is that even a fraction? 9x - 2 = 3? Add 2 and divide by?

OpenStudy (sleepyjess):

Okay. Let's work with what you know. Add 2

Zeronknight (zeronknight):

2 is negative so that cancels out. 3+2=5 9+2=11

OpenStudy (sleepyjess):

That gives \(-\dfrac x9 > 5\). Now, we added to get rid of subtraction. You know we divide to get rid of multiplication. So what do you think we do to get rid of division?

Zeronknight (zeronknight):

Multiply by 9?

OpenStudy (sleepyjess):

Almost, look at the sign, 9 is negative, so multiply by -9

Zeronknight (zeronknight):

ohh. If we multipled 9 by a positive, would the outcome be negative or posiitve?

OpenStudy (sleepyjess):

If we do \(-\dfrac x9 * -\dfrac 91\), the -9 in the denominator of the first fraction cancels with the -9 in the numerator of the second fraction

Zeronknight (zeronknight):

Oh so there is no longer a 9. just X

OpenStudy (sleepyjess):

Correct, now 5*-9 =?

Zeronknight (zeronknight):

45.

Zeronknight (zeronknight):

-45*

Zeronknight (zeronknight):

If we multiply negative by a positive we get negative, right?

OpenStudy (sleepyjess):

Correct, now, when we have an inequality and we have to multiply or divide by a NEGATIVE number, in this case -9, we have to flip the inequality sign. So \(x>-45\) actually becomes \(x < -45\)